# Superpotentials, flat bands and the role of Quantum Geometry for the superfluid stiffness

### v2

@@ Abstract
Here, we demonstrate that contrary to this conventional view, a sufficiently strong superpotential, can indeed generate a flat-band system with a markedly enhanced superconducting transition temperature, $T_{c}$. Probing the superfluid stiffness reveals that, while quantum geometry crucially contributes, its dominant component arises from the residual band dispersion of the occupied states below the Fermi level. Thus, even when the band at the Fermi energy becomes flat, the total kinetic energy remains finite, inherited from lower minibands. This mechanism—absent in twisted moiré systems—establishes a new route to robust superconductivity in superpotential-driven flat-band materials.
Figure 1: Color surface plot of charge density (a) and local order parameter (b) distribution for a checkerboard lattice of $4\times 4$ plaquettes. (c) Band structure for a diagonal cut across the reduced Brillouin zone. (d) Comparison of $T_{c}$ between homogeneous system ($V_{0}/t=0$, blue) and in the presence of a superpotential with $V_{0}/t=2$ (green) and $V_{0}/t=4$ (red). (e) Ratio between $T_{c}$ with a finite superpotential and $T_{c}$ for the homogeneous system. The solid red line in (e) is a fit to $T_{c}/T_{c}^{hom}=\alpha|U|e^{\beta t/|U|}$ with $\alpha=0.0436$ and $\beta=6.663$ as described in the text. Charge concentration: $p=0.5$.
@@ I Electronic multiband structure and critical temperature
which contains a local attraction |U|-|U| between charge fluctuations and we restrict to hopping between nearest neighbor sites ${\bf R}_{i}$, ${\bf R}_{j}$ indicated by $\langle ij\rangle$. For the local potential $V_{i}$ we consider a checkerboard geometry with alternating $4\times 4$ plaquettes within which $V_{i}$ varies between the values $V_{i}=0$ and $V_{i}=-V_{0}$. Eq. (1) is solved via a standard BCS mean-field decoupling, see Supplementary Section I for details. Charge density and order parameter are enhanced on the plaquettes with $V_{i}=-V_{0}$ (Fig. 1a) and the resulting minibands display a flat character for sizeable $V_{0}$, Fig. 1c. The corresponding increased density of states induces a substantial enhancement of $T_{c}$, in particular at weak coupling, Fig. 1d,e. In fact, for flat bands $T_{c}\sim|U|$ [25, 26, 27, 28] while for dispersive bands the standard BCS relation $T_{c}\sim exp(-1/(N(0){|U|}))$ applies, where $N(0)$ denotes the density of states at the Fermi level.
@@ II Superfluid stiffness
defined in the limit $q_{\parallel}=0,q_{\perp}\to 0,\omega=0$ corresponding to the Meissner effect. It is composed of a diamagnetic part, in the present model equivalent to the negative kinetic energy $\langle T_{\alpha}\rangle$ (that for parabolic dispersions equals the conventional $n_{s}/m$), and a paramagnetic term, which arises in inhomogeneous systems. Within our model this term contributes due to the presence of the superpotential scattering. Here we need to point out that contrary to conventional flat band systems we start from a model which has a finite kinetic energy. The multiband structure arises from the folding of the original band and splitting of the minibands due to the superpotential. Thus, even if the kinetic energy is reduced by the superpotential, it generally will remain finite, even if the miniband which crosses $E_{F}$ becomes completely flat.
We also explore a parameter’s regime, where the interaction $|U|$ becomes larger than the hopping $t$. In this regime one cannot neglect the contribution of phase and charge fluctuations to the superfluid stiffness. These fluctuations take into account the phase relaxation of the SC order parameter in response to the applied vector potential and renormalize the stiffness for the inhomogeneous system [29].
Figure 2: Gauge invariant SC current distribution for a vector potential applied along the x-direction. The local gaps are shown as grey background. $U/t=2$, $V_{0}/t=4$, Charge concentration $n_{el}=0.5$.
@@
For the homogeneous system one observes a slight reduction of the stiffness with increasing attractive interaction. Since in this case $D_{s}$ is only due to the diamagnetic part, the behavior of the stiffness is solely determined by the increase of the kinetic energy (i.e. becoming less negative) with increasing order parameter which is the usual behavior for a BCS superconductor [31]. Upon increasing the superpotential $V_{0}$ the overall stiffness gets suppressed with respect to the homogeneous case due to the increasing paramagnetic contribution. In particular, we find that for $V_{0}/t=4$ and $|U|/t<0.5$ the stiffness almost vanishes ($D_{s}\approx 0$) which also holds for the charge stiffness ($D_{c}\approx 0$). For smaller values of $V_{0}/t$ one still finds a finite stiffnes in the limit $|U|/t\to 0$, see Supplementary Figure 3, while the enhancement of $T_{c}$ is much less pronounced, see Fig. 1e (green curve). In case $V_{0}/t=4$ it is therefore remarkable that upon increasing the interaction the stiffness partially recovers and for $|U|/t=3$ approaches a value of $D_{s}/t\approx 0.1$ (red full line) which is almost $17\%$ of the value for the homogeneous system (blue line) so that $\pi D_{s}$ is of the order of $kT_{c}$, cf. Fig. 1d. Note also that the inclusion of phase relaxation via the BCS+RPA approach induces a significant reduction of the paramagnetic contribution to the stiffness from its BCS value (light red curve) for $|U|/t\gtrsim$ 1. While the interaction dependence of $D_{s}$ for $V_{0}/t=4$ mimics the behavior one would expect from the geometric contribution in a flat band SC [11, 13], here the recovery of $D_{s}$ is based on a different mechanism that, however, also includes a positive geometrical contribution. First, let us note that the kinetic term $\langle T_{\alpha}\rangle$ (green line) stays finite due to the contribution of all occupied minibands as explained before. Second, the negative paramagnetic contribution (dashed-dotted line), which for $V_{0}/t=4$ compensates the diamagnetic one at small $|U|/t$, becomes less negative upon increasing the attractive interaction. This can be understood based on the transfer of low energy spectral weight of the optical conductivity to its $\omega=0$ pole with increasing SC gap as it is found in diffusive systems, see Supplementary Section VII. Nevertheless, we will show in the following that a certain geometric contribution is involved in this transfer of spectral weight.
@@
Figure 3: Interaction dependence of the stiffness $D_{s,tot}$for superpotential values $V_{0}/t=0$ (blue) and $V_{0}/t=4$ (red, squares) separated into the diamganetic (mostly intraband) ($D_{s}^{dia}$) and paramagnetic (mostly interband) ($D_{s}^{para}$) contribution. Also shown is the geometric contribution for $r\neq m$ (black) which positively contributes to the stiffness. Lines in the foreground (full symbols) and in the background (open symbols) refer to the RPA (BCS) result, respectively. For $|U|/t=0$ the charge stiffness $D_{c}$ is shown by full circles while $D_{s}$ vanishes in this limit. charge concentration: $p=0.5$.
@@
Figure 4: Color surface plot of ${\cal D}_{r,m}$, see Eq. (5), divided into intraband (a) and interband (b) contributions. charge concentration $p=0.5$; $|U|/t=2$.
@@ III Consideration of disorder
Within our ’mean-field’ BCS approach the critical temperature $kT_{c}$ can be associated with the pairing energy scale whereas $\pi D_{s}$ sets the scale for phase coherence. In the weak-coupling BCS limit one therefore always finds $\pi D_{s}/kT_{c}\gg 1$ since phase coherence is already established when Cooper pairs start to form. In the low-disorder regime where the mean free path is much longer than the inverse Fermi wave vector, $k_{F}l\gg 1$, the stiffness slightly decreases due to the paramagnetic contribution while Anderson’s theorem [32] for the pairing of time-reversed exact eigenstates leaves $T_{c}$ unchanged. The situation changes when $k_{F}l\approx 1$ and Cooper pairs can start to form above the temperature where phase coherence sets in, cf. e.g. Ref. [33], Fig. 3. In this regime the stiffness becomes comparable with $kT_{c}$ eventually leading to a superconductor-insulator transition ($D_{s}=kT_{c}=0$) when disorder exceeds a critical value as e.g. observed in $TiN$, $InO_{x}$ and $NbN$ films [34, 35, 36].
@@
Figure 5: Ratio between the superfluid stiffness and $T_{c}$ as a function of the interaction. Blue and violet curves are for the system without superpotential and impurity potentials $I_{0}/t=0.$ and $I_{0}/t=0.5$, respectively. In the presence of a superpotential ($V_{0}/t=4$, patch size $4\times 4$ the results are shown for impurity strength $I_{0}/t=0$ (black), $I_{0}/t=0.5$ (red), $I_{0}/t=1$ (green). Lines in the foreground (full symbols) and background (open symbols) refer to the RPA (BCS) result, respectively. In case of $V_{0}=I_{0}=0$ RPA and BCS result are identical, whereas for $V_{0}=0$, $I_{0}=0.5t$ the difference is within the size of the symbols. Charge density $n_{el}=0.5$.
We consider two values for the impurity potential $I_{0}/t=0.5$ and $I_{0}/t=1$ which, upon solving the temperature dependent gap equation, have only minor influence on $T_{c}$. The relaxation time is obtained by fitting $\sigma(\omega)$ to the Drude formula, see Supplementary Section V and $k_{F}$ is determined from the average of the Fermi momentum over the Fermi surface of the homogeneous system. For $I_{0}/t=0.5$ this yields a value of $k_{F}l\approx 40$ and a mean free path of $l\approx 23$ in units of the lattice spacing and larger than the superlattice periodicity. In case of $I_{0}/t=1$ one finds $k_{F}l\approx 10$ and $l\approx 6$ which is of the order of the superlattice periodicity (we recall that for the $4\times 4$ plaquette structure the translation vectors are $\vec{a}_{1}=(8,0)$ and $\vec{a}_{2}=(4,4)$ in units of the lattice constant).
@@ IV DISCUSSION
The periodic superpotential can substantially enhance the critical temperature while preserving a sizeable stiffness that scales with the interaction strength. Unlike conventional flat-band systems, where filled bands do not contribute to the kinetic energy and stiffness arises solely from the positive quantum geometric term, our periodically modulated system retains contributions from all minibands below the chemical potential. As a result, even when the top band is flat, a finite diamagnetic response persists. The interband term is overall negative but includes a positive quantum geometric component. Thus, superlattice engineering offers a promising route to enhance Tc, as the system effectively inherits the stiffness of the homogeneous case. It is interesting that recent theoretical work succeeded to predict the topological properties of materials subjected to a superlattice potential [37]. In particular, the periodic patterning of a 2D electron gas [38] reveals a non-zero local Berry phase for antisymmetric triangular superpotentials. This geometry therefore could further sustain the superfluid stiffness because of topology. Consideration of such structures will be considered in future studies within the Usadel equations which allow for the treatment of superpotentials which are incommensurate with the host lattice.
@@
It should be noted that the model here differs from previous work which studied the competition between charge-density wave and SC order parameters [39, 40]. Here instead, we are not considering competing instabilities since the superpotential is imposed on the system via external periodic gating [41] or nanostructuring [42] so that the Fermi level can in principle be tuned into one of the flat bands. This leads to an enhancement of $T_{c}$ with respect to the homogeneous system which at weak coupling can reach one or two orders of magnitude, cf. inset to Fig. 1b. It is to be expected that the influence of phase fluctuations will affect this amplification if the mean free path in the disordered system is of the same order of magnitude as the periodicity of the superlattice which corresponds to the limit $\pi D_{s}/kT_{c}\approx 1$. For strongly disordered NbN films [33] it has been shown that the transition temperature toward a coherent SC state is reduced by a factor $\sim 3$ when this ratio decreases from $\pi D_{s}/kT_{c}\approx 5$ to $\pi D_{s}/kT_{c}\approx 1$ so that our results are still compatible with a significant enhancement of $T_{c}$.
@@
Yet, there are still a few hurdles to overcome in order to realize such a system with increased transition temperature and robust stiffness. First, secondary instabilities such as magnetism may occur when the Fermi level is located in one of the flat bands. The concomitant opening of additional gaps will then compete with the Tc enhancement although the system may remain in a state with sizeable stiffness. Second, as we have seen the system should be sufficiently clean so that the stiffness still dominates the energy scale of the transition temperature. In principle, this requirement should be achievable with LAO/STO 2DEG interfaces where tunable superconductivity has been observed below $0.3$ K [43]. Here one can reach mobilities as high as $10.000$ $cm^{2}/Vs$, corresponding to a mean free path of about $100$ nm [44]. This is well beyond the $40$ nm pitch limit of standard e-beam lithography which defines the possible lower bound periodicity of the metal gate superpotential [45].
@@
To conclude this work demonstrates that applying a superpotential can provide a solid foundation for enhancing superconductivity, even in single-band systems. It thereby paves the way for engineering diverse superpotential architectures as a powerful strategy to design and control superconductivity.

### v3

@@ Abstract
Figure 1: Realization and electronic properties of a device with periodidic gating. (a) Schematic of a 2D superconducting electron gas modulated by a periodic gate potential. A representative platform for realizing such a system is the LaAlO3/SrTiO3 interface. (b) Color surface plot of the modulated charge density for a checkerboard lattice of $4\times 4$ plaquettes. (c) Band structure for a diagonal cut across the reduced Brillouin zone. (d) Color surface plot of the modulated order parameter. Parameters: $U/t=1.5$, $V_{0}/t=4(c)$, $V_{0}/t=2$ (b,d), charge concentration $p=0.5$.
@@
Here, we demonstrate that contrary to this conventional view, a sufficiently strong superpotential, can indeed generate a flat-band system with a markedly enhanced superconducting transition temperature, $T_{c}$. Probing the superfluid stiffness reveals that, while quantum geometry crucially contributes, its dominant component arises from the residual band dispersion of the occupied states below the Fermi level. Thus, even when the band at the Fermi energy becomes flat, the total kinetic energy remains finite, inherited from lower minibands. This mechanism—absent in twisted moiré systems—establishes a new route to robust superconductivity in superpotential-driven flat-band materials and can readily be tested in two-dimensional superconductors as illustrated in Fig. 1a.
@@ I Electronic multiband structure and critical temperature
which contains a local attraction |U|-|U| between charge fluctuations and we restrict to hopping between nearest neighbor sites ${\bf R}_{i}$, ${\bf R}_{j}$ indicated by $\langle ij\rangle$. For the local potential $V_{i}$ we consider a checkerboard geometry with alternating $4\times 4$ plaquettes within which $V_{i}$ varies between the values $V_{i}=0$ and $V_{i}=-V_{0}$. Eq. (1) is solved via a standard BCS mean-field decoupling, see Supplementary Section I for details. Charge density and order parameter are enhanced on the plaquettes with $V_{i}=-V_{0}$ (Fig. 1b and d) and the resulting minibands display a flat character for sizeable $V_{0}$, Fig. 1c. The corresponding increased density of states induces a substantial enhancement of $T_{c}$, in particular at weak coupling, Fig. 2. In fact, for flat bands $T_{c}\sim|U|$ [25, 26, 27, 28] while for dispersive bands the standard BCS relation $T_{c}\sim exp(-1/(N(0){|U|}))$ applies, where $N(0)$ denotes the density of states at the Fermi level.
Figure 2: Critical temperature in the presence of a superpotential. Comparison of $T_{c}$ between homogeneous system ($V_{0}/t=0$, blue) and in the presence of a superpotential with $V_{0}/t=2$ (green) and $V_{0}/t=4$ (red). Inset: Ratio between $T_{c}$ with a finite superpotential and $T_{c}$ for the homogeneous system. The solid red line in (e) is a fit to $T_{c}/T_{c}^{hom}=\alpha|U|e^{\beta t/|U|}$ with $\alpha=0.0436$ and $\beta=6.663$ as described in the text. Charge concentration: $p=0.5$.
@@ II Superfluid stiffness
defined in the momentum limit $q_{\parallel}=0,q_{\perp}\to 0,\omega=0$ corresponding to the Meissner effect. It is composed of a diamagnetic part, in the present model equivalent to the negative kinetic energy $\langle T_{\alpha}\rangle$ (that for parabolic dispersions equals the conventional $n_{s}/m$), and a paramagnetic term, which arises in inhomogeneous systems. Within our model this term contributes due to the presence of the superpotential scattering. Here we need to point out that contrary to conventional flat band systems we start from a model which has a finite kinetic energy. The multiband structure arises from the folding of the original band and splitting of the minibands due to the superpotential. Thus, even if the kinetic energy is reduced by the superpotential, it generally will remain finite, even if the miniband which crosses $E_{F}$ becomes completely flat.
We also explore a parameter regime, where the interaction $|U|$ becomes larger than the hopping $t$. In this regime one cannot neglect the contribution of phase and charge fluctuations to the superfluid stiffness. These fluctuations take into account the phase relaxation of the SC order parameter in response to the applied vector potential and renormalize the stiffness for the inhomogeneous system [29].
Figure 3: Current flow through the periodically modulated superconductor. Illustration of the gauge invariant SC current distribution (white arrows) for a vector potential applied along the x-direction. The local gaps are shown as colored background. $U/t=2$, $V_{0}/t=4$, Charge concentration $n_{el}=0.5$.
@@
For the homogeneous system one observes a slight reduction of the stiffness with increasing attractive interaction. Since in this case $D_{s}$ is only due to the diamagnetic part, the behavior of the stiffness is solely determined by the increase of the kinetic energy (i.e. becoming less negative) with increasing order parameter which is the usual behavior for a BCS superconductor [31]. Upon increasing the superpotential $V_{0}$ the overall stiffness gets suppressed with respect to the homogeneous case due to the increasing paramagnetic contribution. In particular, we find that for $V_{0}/t=4$ and $|U|/t<0.5$ the stiffness almost vanishes ($D_{s}\approx 0$) which also holds for the charge stiffness ($D_{c}\approx 0$). For smaller values of $V_{0}/t$ one still finds a finite stiffness in the limit $|U|/t\to 0$, see Supplementary Figure 3, while the enhancement of $T_{c}$ is much less pronounced, see Fig. 2 (green curve). In case $V_{0}/t=4$ it is therefore remarkable that upon increasing the interaction the stiffness partially recovers and for $|U|/t=3$ approaches a value of $D_{s}/t\approx 0.1$ (red full line) which is almost $17\%$ of the value for the homogeneous system (blue line) so that $\pi D_{s}$ is of the order of $kT_{c}$, cf. Fig. 2. Note also that the inclusion of phase relaxation via the BCS+RPA approach induces a significant reduction of the paramagnetic contribution to the stiffness from its BCS value (light red curve) for $|U|/t\gtrsim$ 1. While the interaction dependence of $D_{s}$ for $V_{0}/t=4$ mimics the behavior one would expect from the geometric contribution in a flat band SC [11, 13], here the recovery of $D_{s}$ is based on a different mechanism that, however, also includes a positive geometrical contribution. First, let us note that the kinetic term $\langle T_{\alpha}\rangle$ (green line) stays finite due to the contribution of all occupied minibands as explained before. Second, the negative paramagnetic contribution (dashed-dotted line), which for $V_{0}/t=4$ compensates the diamagnetic one at small $|U|/t$, becomes less negative upon increasing the attractive interaction. This can be understood based on the transfer of low energy spectral weight of the optical conductivity to its $\omega=0$ pole with increasing SC gap as it is found in diffusive systems, see Supplementary Section VII. Nevertheless, we will show in the following that a certain geometric contribution is involved in this transfer of spectral weight.
@@
Figure 4: Impact of superpotential and gauge invariance on the superfluid stiffness. Interaction dependence of the stiffness $D_{s,tot}$for superpotential values $V_{0}/t=0$ (blue) and $V_{0}/t=4$ (red, squares) separated into the diamganetic (mostly intraband) ($D_{s}^{dia}$) and paramagnetic (mostly interband) ($D_{s}^{para}$) contribution. Also shown is the geometric contribution for $r\neq m$ (black) which positively contributes to the stiffness. Lines in the foreground (full symbols) and in the background (open symbols) refer to the RPA (BCS) result, respectively. For $|U|/t=0$ the charge stiffness $D_{c}$ is shown by full circles while $D_{s}$ vanishes in this limit. charge concentration: $p=0.5$.
@@
Figure 5: Intra- and interband contributions to the superfluid stiffness. Color surface plot of ${\cal D}_{r,m}$, see Eq. (5), divided into intraband (a) and interband (b) contributions. charge concentration $p=0.5$; $|U|/t=2$.
@@ III Consideration of disorder
Within our ’mean-field’ BCS approach the critical temperature $kT_{c}$ can be associated with the pairing energy scale whereas $\pi D_{s}$ sets the scale for phase coherence. In the weak-coupling BCS limit one therefore always finds $\pi D_{s}/kT_{c}\gg 1$ since phase coherence is already established when Cooper pairs start to form. In the low-disorder regime where the mean free path is much longer than the inverse Fermi wave vector, $k_{F}l\gg 1$, the stiffness slightly decreases due to the paramagnetic contribution while Anderson’s theorem [32] for the pairing of time-reversed exact eigenstates leaves $T_{c}$ unchanged. The situation changes when $k_{F}l\approx 1$ and Cooper pairs can start to form above the temperature where phase coherence sets in, cf. e.g. Ref. [33]. In this regime the stiffness becomes comparable with $kT_{c}$ eventually leading to a superconductor-insulator transition ($D_{s}=kT_{c}=0$) when disorder exceeds a critical value as e.g. observed in $TiN$, $InO_{x}$ and $NbN$ films [34, 35, 36].
@@
Figure 6: Ratio between the superfluid stiffness and $T_{c}$. Blue and violet curves are for the system without superpotential and impurity potentials $I_{0}/t=0.$ and $I_{0}/t=0.5$, respectively. In the presence of a superpotential ($V_{0}/t=4$, patch size $4\times 4$) the results are shown for impurity strength $I_{0}/t=0$ (black), $I_{0}/t=0.5$ (red), $I_{0}/t=1$ (green). Lines in the foreground (full symbols) and background (open symbols) refer to the RPA (BCS) result, respectively. In case of $V_{0}=I_{0}=0$ RPA and BCS result are identical, whereas for $V_{0}=0$, $I_{0}=0.5t$ the difference is within the size of the symbols. Charge density $n_{el}=0.5$.
We consider two values for the impurity potential $I_{0}/t=0.5$ and $I_{0}/t=1$ which, upon solving the temperature dependent gap equation, have only minor influence on $T_{c}$. The relaxation time is obtained by fitting $\sigma(\omega)$ to the Drude formula, see Supplementary Section V and $k_{F}$ is determined from the average of the Fermi momentum over the Fermi surface of the homogeneous system. For $I_{0}/t=0.5$ this yields a value of $k_{F}l\approx 40$ and a mean free path of $l\approx 23$ in units of the lattice spacing which is larger than the superlattice periodicity. In case of $I_{0}/t=1$ one finds $k_{F}l\approx 10$ and $l\approx 6$ which is of the order of the superlattice periodicity (we recall that for the $4\times 4$ plaquette structure the translation vectors are $\vec{a}_{1}=(8,0)$ and $\vec{a}_{2}=(4,4)$ in units of the lattice constant).
@@ IV DISCUSSION
The periodic superpotential can substantially enhance the critical temperature while preserving a sizeable stiffness that scales with the interaction strength. Unlike conventional flat-band systems, where filled bands do not contribute to the kinetic energy and a finite stiffness arises solely from the positive quantum geometric term, our periodically modulated system retains contributions from all minibands below the chemical potential. As a result, even when the top band is flat, a finite diamagnetic response persists. The interband term is overall negative but includes a positive quantum geometric component. Thus, superlattice engineering offers a promising route to enhance Tc, as the system effectively inherits the stiffness of the homogeneous case. It is interesting that recent theoretical work succeeded to predict the topological properties of materials subject to a superlattice potential [37]. In particular, the periodic patterning of a 2D electron gas [38] reveals a non-zero local Berry phase for antisymmetric triangular superpotentials. This geometry therefore could further sustain the superfluid stiffness because of topology. Consideration of such structures will be considered in future studies within the Usadel equations which allow for the treatment of superpotentials which are incommensurate with the host lattice.
@@
It should be noted that the model here differs from previous work which studied the competition between charge-density wave and SC order parameters [39, 40]. Here instead, we are not considering competing instabilities since the superpotential is imposed on the system via external periodic gating [41] or nanostructuring [42] so that the Fermi level can in principle be tuned into one of the flat bands. This leads to an enhancement of $T_{c}$ with respect to the homogeneous system which at weak coupling can reach one or two orders of magnitude, cf. inset to Fig. 2. It is to be expected that the influence of phase fluctuations will affect this amplification if the mean free path in the disordered system is of the same order of magnitude as the periodicity of the superlattice which corresponds to the limit $\pi D_{s}/kT_{c}\approx 1$. For strongly disordered NbN films [33] it has been shown that the transition temperature toward a coherent SC state is reduced by a factor $\sim 3$ when this ratio decreases from $\pi D_{s}/kT_{c}\approx 5$ to $\pi D_{s}/kT_{c}\approx 1$ so that our results are still compatible with a significant enhancement of $T_{c}$.
Yet, there are still a few hurdles to overcome in order to realize such a system with increased transition temperature and robust stiffness. First, secondary instabilities such as magnetism may occur when the Fermi level is located in one of the flat bands.
@@
The concomitant opening of additional gaps will then compete with the Tc enhancement although the system may remain in a state with sizeable stiffness. Second, as we have seen the system should be sufficiently clean so that the stiffness still dominates the energy scale of the transition temperature. In principle, this requirement should be achievable with LaAlO3/SrTiO3 interfaces where tunable superconductivity has been observed below $0.3$ K [43]. Here one can reach mobilities as high as $10.000$ $cm^{2}/Vs$, corresponding to a mean free path of about $100$ nm [44]. This is well beyond the $40$ nm pitch limit of standard e-beam lithography which defines the possible lower bound periodicity of the metal gate superpotential [45].
@@
To conclude, this work demonstrates that applying a superpotential can provide a solid foundation for enhancing superconductivity, even in single-band systems. It thereby paves the way for engineering diverse superpotential architectures as a powerful strategy to design and control superconductivity.